# Let Me Think! A Long Chain-of-Thought Can Be Worth Exponentially Many Short Ones

### v1

@@ Abstract
Inference-time computation has emerged as a promising scaling axis for improving large language model reasoning. However, despite yielding impressive performance, the optimal allocation of inference-time computation remains poorly understood. A central question is whether to prioritize sequential scaling (e.g., longer chains of thought) or parallel scaling (e.g., majority voting across multiple short chains of thought). In this work, we seek to illuminate the landscape of test-time scaling by demonstrating the existence of reasoning settings where sequential scaling offers an exponential advantage over parallel scaling. These settings are based on graph connectivity problems in challenging distributions of graphs. We validate our theoretical findings with comprehensive experiments across a range of language models, including models trained from scratch for graph connectivity with different chain of thought strategies as well as large reasoning models. [^1] Code is available here: https://github.com/seyedparsa/let-me-think
@@ 1 Introduction
Large Language Model (LLM) scaling has recently undergone a paradigm shift toward increasing the amount of compute used during inference (SLXK, 25; WSL+, 24), moving beyond traditional axes such as model size, training data and pretraining compute (RNS+, 18; KMH+, 20). Scaling inference-time compute is particularly important for reasoning tasks, and is a key ingredient in OpenAI’s o-series models Ope (24), DeepSeek-R1 GYZ+ (25) among other frontier models MYS+ (25); TDG+ (25); YYZ+ (24); Tea (25).
Despite the impressive performance of these systems, the central question of how to optimally allocate inference-time compute is not yet settled. The main challenge is that the space of strategies that use compute at test time is large and diverse: a wide variety of methods exist WSL+ (24); RR (24); WBF+ (24), each with its own empirical scaling law (WSL+, 24; SLXK, 25). Additionally, different methods can sometimes be combined, which further complicates any analysis.
In this paper, we seek fundamental and general principles that help clarify the landscape of inference-time compute. Since there is a large range of inference-time methods, in order to make progress we categorize methods into two classes (MYS+, 25): (1) parallel scaling and (2) sequential scaling. We review these notions below.
@@
    Parallel scaling refers to generating multiple independent responses in parallel, and aggregating them in some way to output the final solution BJE+ (24); WWS+ (23). The most common aggregation technique is “best-of-$n$”, where a reward function (e.g. another language model or a dedicated verifier (BJE+, 24)) selects the single highest-scoring response as the output. Another widely used aggregation method is majority voting, which determines the final response by choosing the most frequent one among all generated responses (BJE+, 24).
@@
    Sequential scaling encompasses all techniques that do not fall under parallel scaling. The flagship method in this category is Chain-of-Thought (CoT) (WWS+, 22; NAGA+, 22; KGR+, 23), in which an LLM first outputs a chain of reasoning tokens, before outputting its final answer. This may be achieved with one of several strategies to induce longer chains of reasoning in LLMs, such as adding a prompt instruction to “think step by step” KGR+ (23), or forcing a longer chain of thought by replacing end-of-text tokens with “Wait” MYS+ (25), or training with reinforcement learning objectives which can automatically induce longer chains of thought GYZ+ (25).
Consensus has yet to be reached on how to balance both types of scaling most effectively. On the one hand, sequential scaling via long chains of thought has demonstrated particular promise for tackling challenging problems, such as mathematics and coding benchmarks  (GCS+, 25; YTN+, 25; YMLW, 25; LCG+, 25; GYZ+, 25; MYS+, 25). On the other hand, the computational cost of inference grows quadratically in the context window for transformer-based architectures VSP+ (17), making sequential scaling more expensive per-token than parallel scaling. This motivates the main question addressed in this work:
Can we quantify the tradeoff between sequential and parallel scaling for reasoning problems?
@@ 1.1 Our Contributions
Our main contribution is to introduce a reasoning task in which sequential scaling can be exponentially more powerful than parallel scaling. Namely, a small decrease in sequential scale necessitates a large increase in parallel scale to achieve the same level of accuracy. This tradeoff is illustrated in Figure 1 for transformer models evaluated on this task.
Figure 1: Model accuracy with different combinations of parallel and sequential scaling on a graph reasoning task. The sequential scale is the length budget for the chain of thought, and the parallel scale is the number of independent chains of thought generated. The left figure reports the performance of a small transformer model trained for this task, where aggregation is with best-of-$n$. The right figure shows the performance of the frontier DeepSeek-R1-Distill-Qwen-32B reasoning model GYZ+ (25), where aggregation is with majority vote. In both cases, there is a regime in which large increases to parallel scaling are required to compensate for a small decrease in sequential scaling; more details in Section 4.
@@
Our reasoning task is a variant of the basic graph connectivity task. Solving it requires determining whether pairs of vertices are connected by stepping through several edges, and thus it serves as a proxy task modeling multi-step reasoning with CoT on more naturalistic data. Details are in Section 2.
Our task is motivated by a growing theoretical literature on the limitations and capabilities of transformers on graph reasoning tasks (XLZ+, 20; SFH+, 24; ABL+, 24; BBK+, 24; KWLS, 25; SHT, 24), which have found that graph connectivity is challenging for bounded-depth transformers. The reason for this is that graph computation appears to be a sequential problem, yet the transformers’ sequential computation is bounded by their depth, as proved in expressivity results (MS24a, ; MS, 23; Chi, 24).
@@
First, we prove that (a) sequentially scaling with one polynomial-length CoT can solve the connectivity problem, but in contrast (b) parallel scaling by aggregating over polynomially-many $O(1)$-length Chains of Thought cannot succeed. The proof of this theorem leverages recent results on the expressivity of transformers  (MS24a, ; Chi, 24), and requires making complexity-theoretic assumptions; see Section 3.1.
Second, in order to obtain a more fine-grained picture of the landscape and understand the performance of chains of thought with greater than constant length, we abstract transformer computation on graph reasoning tasks with a “Vertex Query Model” of computation. This model of computation is inspired by known limitations of transformers for multi-hop reasoning SHT (24). The Vertex Query Model has the benefit that it is tractable to analyze. Thus, we use it to guide the construction of challenging distributions of “two-path” graphs and “bridge” graphs, for which we give evidence that there is an exponential gap between the performance of sequential and parallel scaling; see Section 3.2.
@@
In Section 4, we empirically validate the challenging distributions of “two-path” and “bridge” graphs motivated by our Vertex Query Model. We use these tasks to test both transformer-based language models trained to solve graph connectivity, as well as leading open-source reasoning models, and we find that there a significant advantage to scaling sequential computation over scaling parallel computation.
In Section 5, we explore training transformers on the graph connectivity problem with reinforcement learning (RL). We observe the emergent behavior that RL training gradually increases the length of the CoT. This behavior mirrors the growth in CoT length that occurs when training DeepSeek-R1 GYZ+ (25) with RL, and supports that the graph connectivity problems studied in this work are a rich enough task to capture many interesting behaviors observed in practice.
@@ 2 Graph reasoning tasks
Motivated by recent work on the expressivity and limitations of constant-depth transformers ABL+ (24); MS24a ; SHT (24); XLZ+ (20); BBK+ (24); KWLS (25), we test models on a graph connectivity task that serves as a basic testbed for reasoning. The most canonical connectivity task that one could consider is $(s,t)$-connectivity, defined below.
Definition 1 ($(s,t)$-connectivity problem).
@@
Our theoretical results and our experiments are for $(s,t_{1},t_{2})$-connectivity in the setting where $G$ consists of two identical, disconnected components, one of the components contains $s$ and $t_{i}$, and the other component contains $t_{3-i}$. The task is inputted as a list of edges, followed by the IDs of $s$, $t_{1}$, and $t_{2}$. See Figure 2 for an example of the input format.
@@ 3.1 Separation Based on Transformer Expressivity Limitations
Theorem 1 (Informal statement of Theorem 4).
@@
In Appendix A we provide the formal statement of the theorem and the full proof of the theorem. For the positive result, the main ingredient is from MS24a , which implies that transformers with polynomial length CoT can implement any polynomial-time algorithm, and therefore can implement breadth-first search which solves the connectivity problem. For the negative result, the expressivity bounds of Chi (24); MS (23) imply that transformers with $O(1)$-length chain-of-thought fall into the class of circuits $\mathsf{TC}^{0}$. Our main insight is that aggregating multiple independently-sampled CoTs is also a $\mathsf{TC}^{0}$ circuit, and therefore is unable to solve $(s,t)$-connectivity under the complexity-theoretic assumption. Finally, we reduce from the $(s,t)$-connectivity problem to the $(s,t_{1},t_{2})$-connectivity problem with a $\mathsf{TC}^{0}$ reduction.
@@ 3.2 Evidence for Separation Based on the Vertex Query Model
This simplifying abstraction is motivated by prior literature. First, constant-depth transformers are known to have limited range for multi-hop reasoning in graphs SHT (24). In chains of thought that output a sequence of nodes, it is reasonable to expect that the next node outputted by a transformer should lie only in a constant-depth neighborhood of the previous nodes or be a randomly-chosen node. These kinds of chains of thought correspond exactly to VQM algorithms. Second, the VQM is closely related to the previously proposed “globality barrier” for transformers learning to reason (Definition 3 of ABL+ (24)). The “globality barrier” suggests that transformers with CoT can only efficiently learn functions such that each CoT step does a local computation – depending on a constant number of entries in the previous chain of thought. The VQM corresponds to such algorithms where the local computations allowed are neighborhood queries.
@@
  • $O(L)$ queries are sufficient: There is a VQM algorithm that executes $L-1$ queries and solves the $(s,t_{1},t_{2})$-connectivity problem with probability 1.
@@
A drawback of the above theorem is that it proves limitations when the number of queries is smaller than the length of the shortest path between $s$ and $\{t_{1},t_{2}\}$. In those situations, it may be impossible for algorithms in the VQM model to certify which $s$ and $t_{i}$ are connected. This raises the question: are there graphs where sequential scale is still beneficial even with more queries than the shortest path length? We provide one such example below, with the “bridge graph” construction. An example of this graph structure is illustrated in Figure 3.
@@
Consider an algorithm in the Restricted Vertex Query Model solving $(s,t_{1},t_{2})$-connectivity on the union of two identical copies of the Bridge$(d,l,2l,0)$ graph, where $s$ is the starting node on one side of the graph, and $t_{1}$ and $t_{2}$ correspond to the copies in the two connected components of the end node of the main path on the other side. For any $\delta\in(0,1)$,
@@
We emphasize that the shortest path between the $s$ and $t_{i}$ is of length $ld$, and the theorem proves that any algorithm in the RVQM has exponentially poor advantage over random guessing even for a number of queries a constant fraction larger than this shortest path. The intuition is that each time the model hits an intersection (that is, a vertex with degree greater than two), it has to guess where to go next, and only has a constant probability per intersection of choosing the shortest path. The proof can be found in Appendix B.
@@ 4 Empirical evidence for benefits of sequential over parallel scaling
Decision: [1]   Path: [15 2 57 27 33 12 41 46 29 54 18 1]
Decision: [1]   Shortest-Path: [15 2 57 27 37 44 29 54 18 1]
@@
4.1 Chain-of-Thought Strategies
@@
The strategy with the shortest CoT that we consider is Shortest-Path, where the training data consists of shortest paths from the source node to the target node. Two other CoT strategies are derived from the trace of a depth-first-search (DFS) starting from the source node and ending at the target node. Path CoT is the path from the source node to the target node in the DFS tree, and DFS CoT is the list of DFS tree nodes ordered by when they are first visited in the DFS trace. The CoT and the final decision are appended to an input prompt to form a training example of the CoT strategy dataset (see Figure 3).
@@ 4.3 Experiment Setup
In our experiments, we let short=$3$, long=$5$, deadend=$3$, and refer to Bridge(d, 3, 5, 3) as Bridge(d). To construct a task of depth $d$, we generate two randomly labeled Bridge(d) graphs, select the first starting node of one of the graphs as the source node $s$, and the last ending nodes of the two graphs as the target nodes $t_{1}$ and $t_{2}$. Finally, to transform the task into a sequence, we list the edges of the graphs in a random order, along with the labels of the source and target nodes as illustrated in Figure 3. For each CoT strategy and Bridge(d) task with depth from 1 to 5, we train a Mistral causal language model (JSM+, 23; WDS+, 20) with 4 hidden layers, 4 attention heads, and intermediate size 128 with a context length of 400 for 200 epochs. In the experiments of each task, we use the same number of training tokens for all CoT strategies, equal to the number of training tokens in 500,000 samples from the Shortest-Path CoT strategy.

### v2

@@ Abstract
Inference-time computation has emerged as a promising scaling axis for improving large language model reasoning. However, despite yielding impressive performance, the optimal allocation of inference-time computation remains poorly understood. A central question is whether to prioritize sequential scaling (e.g., longer chains of thought) or parallel scaling (e.g., majority voting across multiple short chains of thought). In this work, we seek to illuminate the landscape of test-time scaling by demonstrating the existence of reasoning settings where sequential scaling offers an exponential advantage over parallel scaling. These settings are based on graph connectivity problems in challenging distributions of graphs. We validate our theoretical findings with comprehensive experiments across a range of language models, including models trained from scratch for graph connectivity with different chain of thought strategies as well as large reasoning models. Our code is available at https://github.com/seyedparsa/let-me-think.
@@ 1 Introduction
Large Language Model (LLM) scaling has recently undergone a paradigm shift toward increasing the amount of compute used during inference (Snell et al., 2025; Wu et al., 2024), moving beyond traditional axes such as model size, training data and pretraining compute (Radford et al., 2018; Kaplan et al., 2020). Scaling inference-time compute is particularly important for reasoning tasks, and is a key ingredient in OpenAI’s o-series models OpenAI (2024), DeepSeek-R1 Guo et al. (2025) among other frontier models Muennighoff et al. (2025); Team et al. (2025); Yang et al. (2024); Team (2025).
Despite the impressive performance of these systems, the central question of how to optimally allocate inference-time compute is not yet settled. The main challenge is that the space of strategies that use compute at test time is large and diverse: a wide variety of methods exist Wu et al. (2024); Rush and Ritter (2024); Welleck et al. (2024), each with its own empirical scaling law (Wu et al., 2024; Snell et al., 2025). Additionally, different methods can sometimes be combined, which further complicates any analysis.
In this paper, we seek fundamental and general principles that help clarify the landscape of inference-time compute. Since there is a large range of inference-time methods, in order to make progress we categorize methods into two classes (Muennighoff et al., 2025): (1) parallel scaling and (2) sequential scaling. We review these notions below.
@@
    Parallel scaling refers to generating multiple independent responses in parallel, and aggregating them in some way to output the final solution Brown et al. (2024); Wang et al. (2023). The most common aggregation technique is “best-of-nn”, where a reward function (e.g. another language model or a dedicated verifier (Brown et al., 2024)) selects the single highest-scoring response as the output. Another widely used aggregation method is majority voting, which determines the final response by choosing the most frequent one among all generated responses (Brown et al., 2024).
@@
    Sequential scaling encompasses all techniques that do not fall under parallel scaling. The flagship method in this category is Chain of Thought (CoT) (Wei et al., 2022; Nye et al., 2022; Kojima et al., 2023), in which an LLM first outputs a chain of reasoning tokens, before outputting its final answer. This may be achieved with one of several strategies to induce longer chains of reasoning in LLMs, such as adding a prompt instruction to “think step by step” Kojima et al. (2023), or forcing a longer chain of thought by replacing end-of-text tokens with “Wait” Muennighoff et al. (2025), or training with reinforcement learning objectives which can automatically induce longer chains of thought Guo et al. (2025).
Consensus has yet to be reached on how to balance both types of scaling most effectively. On the one hand, sequential scaling via long chains of thought has demonstrated particular promise for tackling challenging problems, such as mathematics and coding benchmarks  (Gandhi et al., 2025; Yeo et al., 2025; Yang et al., 2025c; Li et al., 2025; Guo et al., 2025; Muennighoff et al., 2025). On the other hand, the computational cost of inference grows quadratically in the context window for transformer-based architectures Vaswani et al. (2017), making sequential scaling more expensive per-token than parallel scaling. This motivates the main question addressed in this work:
Can we quantify the trade-off between sequential and parallel scaling for reasoning problems?
@@ 1.1 Our contributions
Our main contribution is to introduce a reasoning task in which sequential scaling can be exponentially more powerful than parallel scaling. Namely, a small decrease in sequential scale necessitates a large increase in parallel scale to achieve the same level of accuracy. This tradeoff is illustrated in Figure˜1 for transformer models evaluated on this task.
Figure 1: Model accuracy with different combinations of parallel and sequential scaling on a graph reasoning task. The sequential scale is the length budget for the chain of thought, and the parallel scale is the number of independent chains of thought generated. The left figure reports the performance of a small transformer model trained for this task, where aggregation is with best-of-$n$. The right figure shows the performance of the frontier DeepSeek-R1-Distill-Qwen-32B reasoning model Guo et al. (2025), where aggregation is with majority vote. In both cases, there is a regime in which large increases to parallel scaling are required to compensate for a small decrease in sequential scaling; more details in Section˜4.
@@
Our reasoning task is a variant of the basic graph connectivity task. Solving it requires determining whether pairs of vertices are connected by stepping through several edges, and thus it serves as a proxy task modeling multi-step reasoning with CoT on more naturalistic data. Details are in Section˜2.
Our task is motivated by a growing theoretical literature on the limitations and capabilities of transformers on graph reasoning tasks (Xu et al., 2020; Sanford et al., 2024a; Abbe et al., 2024; Besta et al., 2024; Kim et al., 2025; Sanford et al., 2024b), which have found that graph connectivity is challenging for bounded-depth transformers. The reason for this is that graph computation appears to be a sequential problem, yet the transformers’ sequential computation is bounded by their depth, as proved in expressivity results (Merrill and Sabharwal, 2024b, 2023; Chiang, 2024).
@@
First, we prove that (a) sequentially scaling with one polynomial-length CoT can solve the connectivity problem, but in contrast (b) parallel scaling by aggregating over polynomially-many $O(1)$-length Chains of Thought cannot succeed. The proof of this theorem leverages recent results on the expressivity of transformers  (Merrill and Sabharwal, 2024b; Chiang, 2024), and requires making complexity-theoretic assumptions; see Section 3.1.
Second, in order to obtain a more fine-grained picture of the landscape and understand the performance of chains of thought with greater than constant length, we abstract transformer computation on graph reasoning tasks with a “Vertex Query Model” of computation. This model of computation is inspired by known limitations of transformers for multi-hop reasoning Sanford et al. (2024b). The Vertex Query Model has the benefit that it is tractable to analyze. Thus, we use it to guide the construction of challenging distributions of “two-path” graphs and “bridge” graphs, for which we give evidence that there is an exponential gap between the performance of sequential and parallel scaling; see Section˜3.2.
@@
In Section˜4, we empirically validate the challenging distributions of “two-path” and “bridge” graphs motivated by our Vertex Query Model. We use these tasks to test transformer-based language models trained to solve graph connectivity, and find that there is a significant advantage to scaling sequential computation over scaling parallel computation. We then extend this empirical investigation to leading open-source reasoning models, evaluating their performance on the graph connectivity task as well as on the more complex AIME2024 of America (2024) dataset. The results reveal a consistent trend favoring sequential scaling over parallel scaling.
In Figure˜7, we explore training transformers on the graph connectivity problem with reinforcement learning (RL). We observe the emergent behavior that RL training gradually increases the length of the CoT. This behavior mirrors the growth in CoT length that occurs when training DeepSeek-R1 Guo et al. (2025) with RL, and supports that the graph connectivity problems studied in this work are a rich enough task to capture many interesting behaviors observed in practice.
@@ 2 Graph reasoning tasks
Motivated by recent work on the expressivity and limitations of constant-depth transformers Abbe et al. (2024); Merrill and Sabharwal (2024b); Sanford et al. (2024b); Xu et al. (2020); Besta et al. (2024); Kim et al. (2025), we test models on a graph connectivity task that serves as a basic testbed for reasoning. The most canonical connectivity task that one could consider is $(s,t)$-connectivity, defined below.
Definition 1 $($s,t)(s,t)-connectivity problem).
@@
Our theoretical results and our experiments are for $(s,t_{1},t_{2})$-connectivity in the setting where $G$ consists of two identical, disconnected components, one of the components contains $s$ and $t_{i}$, and the other component contains $t_{3-i}$. The task is inputted as a list of edges, followed by the IDs of $s$, $t_{1}$, and $t_{2}$. See Figure˜2 for an example of the input format.
@@ 3.1 Separation based on transformer expressivity limitations
Theorem 1 (Informal statement of Theorem˜4).
@@
In Appendix A we provide the formal statement of the theorem and the full proof of the theorem. For the positive result, the main ingredient is from Merrill and Sabharwal (2024b), which implies that transformers with polynomial length CoT can implement any polynomial-time algorithm, and therefore can implement breadth-first search which solves the connectivity problem. For the negative result, the expressivity bounds of Chiang (2024); Merrill and Sabharwal (2023) imply that transformers with $O(1)$-length chain-of-thought fall into the class of circuits $\mathsf{TC}^{0}$. Our main insight is that aggregating multiple independently-sampled CoTs is also a $\mathsf{TC}^{0}$ circuit, and therefore is unable to solve $(s,t)$-connectivity under the complexity-theoretic assumption. Finally, we reduce from the $(s,t)$-connectivity problem to the $(s,t_{1},t_{2})$-connectivity problem with a $\mathsf{TC}^{0}$ reduction.
@@ 3.2 Evidence for separation based on the vertex query model
This simplifying abstraction is motivated by prior literature. First, constant-depth transformers are known to have limited range for multi-hop reasoning in graphs Sanford et al. (2024b). In chains of thought that output a sequence of nodes, it is reasonable to expect that the next node outputted by a transformer should lie only in a constant-depth neighborhood of the previous nodes or be a randomly-chosen node. These kinds of chains of thought correspond exactly to VQM algorithms. Second, the VQM is closely related to the previously proposed “globality barrier” for transformers learning to reason (Definition 3 of Abbe et al. (2024)). The “globality barrier” suggests that transformers with CoT can only efficiently learn functions such that each CoT step does a local computation – depending on a constant number of entries in the previous chain of thought. The VQM corresponds to such algorithms where the local computations allowed are neighborhood queries.
@@
  • $O(L)$ queries are sufficient: There is a VQM algorithm that executes L1L-1 queries and solves the $(s,t_{1},t_{2})$-connectivity problem with probability 1.
@@
A drawback of the above theorem is that it proves limitations when the number of queries is smaller than the length of the shortest path between $s$ and $\{t_{1},t_{2}\}$. In those situations, it may be impossible for algorithms in the VQM model to certify which $s$ and $t_{i}$ are connected. This raises the question: are there graphs where sequential scale is still beneficial even with more queries than the shortest path length? We provide one such example below, with the “bridge graph” construction. An example of this graph structure is illustrated in Figure˜3.
@@
Consider an algorithm in the Restricted Vertex Query Model solving $(s,t_{1},t_{2})$-connectivity on the union of two identical copies of the Bridge(d,l,2l,0)(d,l,2l,0) graph, where $s$ is the starting node on one side of the graph, and $t_{1}$ and $t_{2}$ correspond to the copies in the two connected components of the end node of the main path on the other side. For any $\delta\in(0,1)$,
@@
We emphasize that the shortest path between the $s$ and $t_{i}$ is of length $ld$, and the theorem proves that any algorithm in the RVQM has exponentially poor advantage over random guessing even for a number of queries a constant fraction larger than this shortest path. The intuition is that each time the model hits an intersection (that is, a vertex with degree greater than two), it has to guess where to go next, and only has a constant probability per intersection of choosing the shortest path. The proof can be found in Appendix˜B.
@@ 4 Empirical evidence for benefits of sequential over parallel scaling
Decision: [1]
  Path: [15 2 57 27 33 12 41 46 29 54 18 1]
Decision: [1]
  Shortest-Path: [15 2 57 27 37 44 29 54 18 1]
@@
4.1 Chain of thought strategies
@@
The strategy with the shortest CoT that we consider is Shortest-Path, where the training data consists of shortest paths from the source node to the target node. Two other CoT strategies are derived from the trace of a depth-first-search (DFS) starting from the source node and ending at the target node. Path CoT is the path from the source node to the target node in the DFS tree, and DFS CoT is the list of DFS tree nodes ordered by when they are first visited in the DFS trace. The CoT and the final decision are appended to an input prompt to form a training example of the CoT strategy dataset (See Figure˜3).
@@ 4.3 Experiment setup
In our experiments, we let short=$3$, long=55, deadend=33, and refer to Bridge(d, 3, 5, 3) as Bridge(d). To construct a task of depth $d$, we generate two randomly labeled Bridge(d) graphs, select the first starting node of one of the graphs as the source node $s$, and the last ending nodes of the two graphs as the target nodes $t_{1}$ and $t_{2}$. Finally, to transform the task into a sequence, we list the edges of the graphs in a random order, along with the labels of the source and target nodes as illustrated in Figure˜3. For each CoT strategy and Bridge(d) task with depth from 1 to 5, we train a Mistral causal language model (Jiang et al., 2023; Wolf et al., 2020) with 4 hidden layers, 4 attention heads, and intermediate size 128 with a context length of 400 for 200 epochs. In the experiments of each task, we use the same number of training tokens for all CoT strategies, equal to the number of training tokens in 500,000 samples from the Shortest-Path CoT strategy.